# UNGT: Ultrasound Nasogastric Tube Dataset for Medical Image Analysis

### v1

@@ 5.2 Classification
We appraise the classification performance among different SOTA methods on the proposed UNGT dataset in table 2. An insightful examination delineates that the GhostNet and CECT showcase superior performance than the remaining SOTA approaches. Specifically, GhostNet demonstrates the optimal PRE and SPE of 89.30% and 80.13%, while CECT presents the highest ACC, SEN, NPV, and FOS of 85.90%, 91.07%, 82.73%, and 89.33%. Conversely, the InceptionNext demonstrates somewhat undesirable results with an ACC of 72.27%. It is worth noting that the standard deviations observed across varying methods are relatively pronounced, which may be ascribed to the constrained size of the introduced UNGT dataset. Even though, several approaches such as RegNetX and MobileNetV4 demonstrate remarkable standard deviations. Taking ACC as an example, these two approaches present standard deviations of 0.53% and 0.33% respectively, demonstrating remarkable training stability under the small data regime. In contrast, MobileViT exhibits a significantly elevated ACC standard deviation of 4.08%, evidencing potential dramatic volatility as training proceeds. The exceptional performance metrics achieved by GhostNet and CECT underscore their effectiveness, while the outstanding standard deviations exhibited by RegNetX and MobileNetV4 emphasize their superior training stability.
@@ 6 Conclusion
In this paper, we develop a novel UNGT dataset to mitigate the dataset gap. The proposed dataset comprises 493 images from 110 patients, featuring an average image resolution of approximately 879 $\times$ 583. Besides, we introduce a semi-supervised AAMS approach to tackle data limitation and imbalance concurrently. We conduct comprehensive segmentation experiments and the results demonstrate that the developed AAMS shows manifest performance leadership compared with prevalent SOTA methods. Moreover, we perform extensive classification studies on varying SOTA approaches and compare their performance. We believe the first US NGT dataset and novel medical segmenter for tackling data limitation and imbalance can provide valuable insights and benefit the broader medical research community.
The potential approach limitations and further research perspectives are twofold. For one, the introduced UNGT dataset comprises a limited number of images and faces severely unbalanced structures. To this end, collecting and annotating supplementary images can be accounted for, especially for the underrepresented tube and pancreas structures. With these, the enlarged dataset can comprise a more balanced category distribution. For another, the proposed AAMS trains varying structures with identical iterations. Accordingly, training infrequent or minor structures with additional iterations than the common ones can be beneficial. For example, objects such as the tube and pancreas could be trained with more iterations than other categories such as the liver and stomach. This strategy involves processing all channels in each batch during standard training while dedicating additional iterations to sparse or small categories to emphasize their features. Such an approach could improve the performance of sparse or small objects without compromising the learning of others. The weighting scheme can then be optimized based on such modifications for further performance enhancement.

### v2

@@ 5.2 Classification
We appraise the classification performance among different SOTA methods on the proposed UNGT dataset in table 2. An insightful examination delineates that the GhostNet and CECT showcase superior performance compared to the remaining SOTA approaches. Specifically, CECT presents the highest ACC, SEN, NPV, and FOS of 85.90%, 91.07%, 82.73%, and 89.33%, and GhostNet demonstrates the optimal PRE and SPE of 89.30% and 80.13%. Conversely, the InceptionNext demonstrates somewhat undesirable results with an ACC of 72.27%. It is worth noting that the standard deviations observed across varying methods are relatively pronounced, which may be ascribed to the constrained size of the introduced UNGT dataset. Despite this, approaches such as MobileNetV4 and RegNetX demonstrate ideal standard deviations. Taking ACC as an example, these two approaches present standard deviations of 0.33% and 0.53%, demonstrating remarkable training stability under the small data regime. In contrast, MobileViT exhibits a significantly elevated ACC standard deviation of 4.08%, evidencing potential dramatic volatility as training proceeds. The exceptional performance metrics achieved by CECT and GhostNet underscore their effectiveness, while the outstanding standard deviations exhibited by MobileNetV4 and RegNetX emphasize their superior training stability.
@@
6 Ablation and Generalization
6.1 Ablation Study
Figure 7: Ablation study of the ADW scheme and MAA block across the tube and pancreas on the introduced UNGT dataset. (a) Tube, and (b) pancreas.
We illustrate the ablation study of the ADW scheme and MAA block across the tube and pancreas on the proposed UNGT dataset in fig. 7. The results indicate that prediction accuracy decreases with the removal of a model component, and removing both components leads to a more severe performance drop. For the second image, the AAMS correctly identifies the absence of the tube structure, while its two variants fail to do so. As for the fifth figure, AAMS consistently captures the tube absence, whereas its variant without the ADW scheme and MAA block fails in this regard. Regarding the eighth image, the AAMS produces the most accurate pancreas shape, while its two ablation variants substantially underestimate the pancreas size. In the eleventh figure, the AAMS delivers the most complete pancreas, whereas both ablation variants output fragmented representations, particularly the variant without both the ADW scheme and MAA block. The performance degradation after removing different numbers of components underscores their importance in the proposed AAMS.
6.2 Generalization Evaluation
Table 3: Segmentation results across AAMS and SOTA approaches on the CAMUS dataset. LVendo, LVepi, and LA present the left ventricle endocardium, epicardium contour, and left atrium.
Model DSC SEN PRE
LVendo LVepi LA LVendo LVepi LA LVendo LVepi LA
UNet [44] 93.730.02 87.700.09 89.080.05 94.080.17 87.910.29 89.330.40 93.820.14 87.910.12 90.380.40
TransUnet [6] 93.430.03 87.550.15 88.930.05 93.200.42 89.110.92 90.350.72 94.200.47 86.510.57 88.970.66
UNext [53] 92.470.02 85.770.05 87.250.12 93.170.01 86.800.41 88.430.01 92.420.01 85.270.30 87.760.22
CMUNeXt [49] 93.250.02 87.150.05 88.760.05 93.840.21 88.390.56 90.441.29 93.180.15 86.390.61 88.551.29
MedT [52] 91.500.06 83.790.03 85.350.25 92.810.46 85.580.07 87.280.81 90.910.34 82.720.02 85.410.31
CMUNet [51] 93.730.01 87.910.03 89.500.10 93.910.16 89.100.02 89.670.36 94.010.19 87.170.04 90.640.49
CCT [39] 92.210.27 83.941.03 88.440.12 90.551.62 82.762.22 89.890.95 94.701.28 85.950.39 88.731.18
CPS [7] 91.741.24 82.912.38 88.820.52 89.352.15 80.033.70 90.600.39 95.060.03 87.050.71 88.530.69
CTCT [34] 92.510.27 84.700.46 89.000.15 90.691.19 83.631.66 90.220.16 95.130.87 86.491.02 89.290.38
ICT [54] 92.890.03 85.960.16 88.920.25 91.730.44 85.460.76 90.150.59 94.770.52 87.110.54 89.230.65
R-Drop [61] 92.270.30 84.700.64 88.590.07 90.100.91 83.941.42 90.250.85 95.360.43 86.231.10 88.501.00
URPC [35] 92.960.11 85.700.58 88.870.14 91.900.43 85.311.93 90.490.62 94.650.34 86.771.00 88.850.46
MGCC [50] 93.790.02 88.070.00 89.380.23 94.060.00 89.030.20 89.960.08 93.980.03 87.550.19 90.250.37
AAMS (Ours) 94.110.00 88.480.00 90.330.00 94.610.01 89.330.20 91.050.04 94.030.00 88.070.17 90.840.08
We present the segmentation results across AAMS and SOTA methods on the CAMUS dataset in table 3. It is evident that our AAMS achieves superior overall performance compared to SOTA methods to different degrees on four structures. Considering the left ventricle endocardium, our AAMS attains the highest DSC and SEN of 94.11% and 94.61%. Although R-Drop yields the optimal PRE of 95.36%, it demonstrates a comparatively low SEN of 90.10%. In contrast, the MedT produces suboptimal results under this setup with a DSC of 91.50%. Regarding the epicardium contour, the AAMS achieves the best DSC, SEN, and PRE of 88.48%, 89.33%, and 88.07%. The lowest performance is observed on CPS with a DSC of 82.91%. For the left atrium, the proposed AAMS presents the optimal DSC, SEN, and PRE of 90.33%, 91.05%, and 90.84%. The MedT consistently presents undesirable results with a DSC of 85.35%. The performance superiority of our AAMS highlights its generalization, even though the advantages of the ADW scheme are not fully exploited due to the comparable structure size.
@@ 7 Conclusion
In this paper, we develop a novel UNGT dataset to mitigate the dataset gap. The proposed dataset comprises 493 images from 110 patients, featuring an average image resolution of approximately 879 $\times$ 583. Images obtained from different US machines introduce heterogeneity and enhance the robustness and real-world applicability of the dataset. Besides, we introduce a semi-supervised AAMS approach to tackle data limitation and imbalance concurrently. The combination of semi‑supervised learning and adaptive class weighting ensures its effectiveness and transferability for data shortage or highly imbalanced tasks, such as rare‑lesion or under‑represented organ segmentation. As a best‑practice guideline, we recommend deploying dynamic inverse‑frequency weights and maintaining an unlabeled pool that covers all classes to maximize overall performance. We conduct comprehensive segmentation experiments, and the results demonstrate that the developed AAMS shows manifest performance leadership compared with prevalent SOTA methods. Moreover, we perform extensive classification studies on varying SOTA approaches and compare their performance. We believe the first US NGT dataset and novel medical segmenter for tackling data limitation and imbalance can provide valuable insights and benefit the broader medical research community.
The potential approach limitations and further research perspectives are twofold. For one, the introduced UNGT dataset comprises a limited number of images and encounters severely unbalanced structures. To this end, collecting and annotating supplementary images can be accounted for, especially for the underrepresented tube and pancreas structures. With these, the enlarged dataset can comprise a more balanced category distribution. For another, the proposed AAMS trains varying structures with identical iterations. Accordingly, training infrequent or minor structures with additional iterations than the common ones can be beneficial. For example, objects like the tube and pancreas could be trained with more iterations than the liver and stomach. This strategy involves processing all channels in each batch during standard training while dedicating additional iterations to sparse or small categories to emphasize their features. Such an approach could improve the performance of sparse or small objects without compromising the learning of others. The weighting scheme can then be optimized based on such modifications for further performance enhancement.